# DQ-DETR: DETR with Dynamic Query for Tiny Object Detection

### v2

@@ 5 Conclusion
In this paper, we analyze that the fixed number and position of queries in previous DETR-like methods are unsuitable for aerial datasets and propose a new end-to-end transformer detector DQ-DETR with categorical counting module, counting-guided feature enhancement, and dynamic query selection. Our DQ-DETR dynamically adjusts the number of object queries used for detection to solve the imbalance number of instances between different aerial images. Also, we improve the positional information of queries, making the decoder easier to localize the tiny object. DQ-DETR is the first DETR-like model focusing on tiny object detection and achieves 30.2% $\mathrm{AP}$, which is the state-of-the-art of AI-TOD-V2. The result shows that our proposed DQ-DETR improves the performance of detecting tiny objects, outperforming all previous CNN-based detectors and DETR-like methods on the AI-TOD-V2 dataset with ResNet50 as the backbone.
6 Supplementary
6.1 Quantitative Result
6.1.1 FP/FN Under Different Density Situation
In table 8, we explore our DQ-DETR’s performance under different density situations, including sparse and dense images. We classify images with less than 100 objects as sparse and images with over 900 objects as dense. LRP FP and LRP FN [30] are used as the evaluation metric. Unlike AP metrics, a lower LRP value implies better performance. Previous DETR-like models apply a fixed number of object queries, while our DQ-DETR uses a dynamic number of queries depending on the object’s density in the picture.
DINO-DETR uses a fixed number of 900 queries for detection, no matter whether in dense or sparse situations. The number of queries exceeds the number of objects in the sparse image and hence introduces many underlying false positive samples (FP). In contrast, for dense images, the number of queries DINO-DETR uses is far less than the number of objects in images, which is beyond the detection capability of DINO-DETR, leaving lots of instances undetected (FN) and causes a large LRP FN score. Our proposed DQ-DETR dynamically adjusts the number of queries used for detection, resulting in fewer FP in sparse images and fewer FN in dense images.
Table 8: LRP FP and LRP FN score under different density situations in AI-TOD-V2. DINO-DETR is compared as the baseline.
Method Situation LRP FP LRP FN
DINO-DETR Sparse 29.4 40.7
Dense 36.8 75.1
DQ-DETR Sparse 25.7 36.4
Dense 35.4 51.5
6.1.2 Ablation of Categorical Counting Module
In categorical counting module, we categorize the the number of objects $N$ per image into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. We selected the numbers 10, 100, and 500 based on the AI-TOD-V2 dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Further, we only classify the number of objects $N$ into four levels due to the long-tail distribution of the training samples. For the $N>500$ situation, there are only 46 training images in this situation, which is much fewer than other cases and leads to a poor classification performance with only 56.6% accuracy. Although there are at most 2267 instances per image in the AI-TOD-V2 dataset, the long-tail distribution of the training samples restricts us from classifying the number of instances $N$ per image in a more detail manner.
Table 9 and Table 10 demonstrates the detection performance and the accuracy of the classification task in our categorical counting module. We can observe that if we classify the the number of objects $N$ into more classes, e.g. 5 classes, AP drops 1.4 compared to 4 classes situation. That is because the poor classification results from the categorical counting module will directly affect the number of object queries used for detection and the inappropriate number of queries might harm the detection performance. For 5 classes classification, while the total classification accuracy maintains 93.8%, the accuracy in the $500<$ N $\leq 900$, and $N>900$ situations are only 37.1% and 57.4%. Since there are only few training images, the categorical counting module doesn’t perform well in these two situations and further impacts the detection performance. Hence, we only categorize the the number of objects $N$ per image into four levels without partition the $N>500$ situation into more detailed settings.
Table 9: Ablation of categorical counting module. DINO-DETR is compared as the baseline.
Method $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
Baseline 25.9 12.7 25.3 32.0 39.7
Classification (4cls) 30.2 15.3 30.5 36.5 44.6
Classification (5cls) 28.8 14.3 29.2 34.1 43.1
Table 10: The classification accuracy of our categorical counting module with different numbers of classes.
#Objects in image Accuracy(%) @ 4cls Accuracy(%) @ 5cls #Sample
N $\leq 10$ 97.7 97.5 8674
$10<$ N $\leq 100$ 90.5 89.3 4393
$100<$ N $\leq 500$ 86.5 83.2 905
$500<$ N $\leq 900$ 56.5 37.1 35
$900<$ N - 54.4 11
Total 94.6 93.8 14018
6.2 The categorical counting module (CCM) is dataset distribution-depend.
Since the characteristics of different datasets may vary a lot, it is vital to tailor our categorical counting module depending on the dataset property and determined how many queries should be used for object detection. In categorical counting module for AI-TOD-V2, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. It is worth noting that the hyperparameters 10, 100, 500 is tailored for AI-TOD-V2. For other datasets, the hyperparameters used in the CCM can be pipelined into a logical process using the mean and variance of the objects per image in the dataset to avoid long-tail distribution. There is no need to hand-design the CCM for different datasets.
Algorithm 1 Categorical Counting Pseudo Code
1:function Categorical-Counting($features$)
2:     $mean\leftarrow\text{Mean}(dataset)$
3:     $var\leftarrow\text{Variance}(dataset)$
4:     $GT\leftarrow\text{Predict}(features)$
5:     $class1\leftarrow\{GT<mean-var\}$
6:     $class2\leftarrow\{mean-var\leq GT<mean\}$
7:     $class3\leftarrow\{mean\leq GT<mean+var\}$
8:     $class4\leftarrow\{GT\geq mean+var\}$
9:end function
6.3 Visualization
Fig. 3 visualizes the detection feature maps of our DQ-DETR compared with Deformable-DETR. We can observe that DQ-DTER generates a more informative feature to detect tiny objects. In comparison, the feature from Deformable-DETR lacks of instances information, which leads to many object undetected (FN). Also, since Deformable-DETR uses a lot of object queries whose positional information is not refined for detecting tiny objects, there are many false positive in the detetcion result.
Figure 3: Visualization of detection results and feature maps. The green, red, and blue boxes represent TP, FP, and FN, respectively.

### v3

@@ 5 Conclusion
In this paper, we analyze that the fixed number and position of queries in previous DETR-like methods are unsuitable for detecting tiny objects in aerial datasets and propose a new end-to-end transformer detector DQ-DETR with a categorical counting module, counting-guided feature enhancement, and dynamic query selection. Our DQ-DETR dynamically adjusts the number of object queries used for detection to solve the imbalance of instances between different aerial images. Also, we improve the positional information of queries, making it easier for the decoder to localize the tiny object. DQ-DETR is the first DETR-like model focusing on tiny object detection and achieves 30.2% $\mathrm{AP}$, which is the state-of-the-art of AI-TOD-V2. The result shows that our proposed DQ-DETR improves the performance of detecting tiny objects, outperforming all previous CNN-based detectors and DETR-like methods on the AI-TOD-V2 dataset with ResNet50 as the backbone.
Acknowledgment
This work is partially supported by the National Science and Technology Council, Taiwan under Grants NSTC-112-2221-E-A49-059-MY3, NSTC-112-2221-E-A49-094-MY3, NSTC-112-2628-E-002-033-MY4 and NSTC-112-2634-F-002-002-MBK, and was financially supported in part by the Center of Data Intelligence: Technologies, Applications, and Systems, National Taiwan University (Grants: 113L900901/113L900902/113L900903), from the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education, Taiwan.